EVE 9/27/99 shift summary

Alexander Ledovskoy (ledovsko@fnal.gov)
Tue, 28 Sep 1999 00:11:02 -0500

Shift Summary for EVE shift, 27-SEP-99

On Shift: Ricardo, Sasha L.

Hours of beam available at all intensities: 5.00h
Hours of beam available at nominal intensity: 5:00h
Hours of E799 data (special runs not included): 3.50h
Hours of special runs (laser scans, muon runs...): 1.50h

--------------------------------------------------------------
Accounting of lost time:

a. Beam time:

0.20h Pipeline resets, KFRENDs.
0.45h Tape change, Absorber insert, Run inits.
0.50h Taking data with tripped NM3S and recovery.
0.35h Trigger modifications.

b. 3.50h were lost on recovery from dropped interlocks, twice.

========================================================================
Runs ended this shift:

14745 BEAM_799 10.2M events (no absorber)
14746 COSMICS
14747 BEAM_799 3.01M events (w/absorber)
14748 LASER_SCAN

Tapes brought to FCC this shift:

UPA037-039, UPB037-039, UPC037-039, UPS013

========================================================================
Accesses during shift:

2 Search+Secure parties.

========================================================================
Overview of shift:

3 parts of smooth running
3 parts of no beam
1 part of lost beam due to KTeV problems.
Mix well, serve chilled.

========================================================================
Description of Shift:


run spill time comment
------------------------------------------------------------------------
14745 15:30 Smooth E799 data taking at 5.5E12 without absorber

239 17:30 NW Quadrant in CsI dropped out of readout.
No ktevana errors. Pipeline reset does not
help. This means drt_fresh_start
Tapes are 85% full. Stop the run

ADGN constants are updated (requested by SET)

14746 Cosmic run for JasonH to measure new gain of
an inefficient Etot channel 2362 after threshold
adjustment.

Beam stop in. BobT moved in the absorber.

14747 1 18:09 BEAM_799 started. We didn't do fresh_start. Init
of the new run was enough to clear CsI errors.

18:15 MCR called. Vacuum window shutter is down, interlocks
are dropped. They are investigating.

19:05 Search'n'Secure is complete.

44 19:06 Beam is back

61 19:30 Noticed that NM3S (final sweeper) tripped off
(when did that happened? spill 22!)
Up to this point data is no good in this run.

65 Taking data again.

73 B11 (Hyperon triger) is modified: HCC_GE1 is changed
to HCC_GE2

19:50 Theo and Ashkan resolved the SDAQ problem
of killing itself every 5 min (SDAQ heartbeat was
generated by regenerator movement! Not anymore.)

78 CsI 1381 has bad capid, gain drift and laser RMS
Pipieline reset. That fixes the problem.

84 Trigger B11 (Hyperon) prescale changed from 20 to 7
Trigger B14 (Accidentals) prescale changed from
30K to 15K

20:45 Interlocks dropped again.
Same door (rollup door or the door next to it)
Experts are on the way. Time to do laser scan.
Run stoped.

14748 21:00 LASER_SCAN started
21:45 Laser scan in finished. Still no beam.

14749 21:55 BEAM_799 started

23:15 Second Search.and.Secure completed

67 23:22 Beam is back
Smooth running untill end of the shift.

========================================================================
Follow up Notes on Previous Shift:

ADGN constants were updated
Absorber-in configuration was re-established.

========================================================================
Items to Follow up on Next Shift:

Page Arun ~30 min before next tape change.